9 French Corporation wishes to hire Leslie as a consultant to design a comprehensive staff training program. The project is expected to take one year, and the parties have agreed to a tentative price of $68,000. French Corporation has proposed payment of one-half of the fee now, with the remainder paid in one year when the project is complete. Use Appendix A and Appendix B. 5 a. If Leslie expects her marginal tax rate to be 25 percent this year and 35 percent next year, calculate the after-tax net present value of this contract to Leslie, using a 6 percent discount rate. b. French Corporation expects its marginal tax rate to be 31 percent both years. Calculate the net present value of French's after-tax cost to enter into this contract using a 6 percent discount rate. c-1. Given that Leslie expects her tax rate to increase next year, she would prefer to receive more of the income from the project up front. Consider an alternative proposal under which French pays Leslie $48,000 this year, and $18,000 in one year when the contract is complete. Calculate the after-tax benefit of this counterproposal to Leslie and the after-tax cost to French. c-2. Are both parties better off under this alternative than under the original plan? points 01:17:24 Skipped eBook Complete this question by entering your answers in the tabs below. Req A Req B Req C1 Req C2 If Leslie expects her marginal tax rate to be 25 percent this year and 35 percent next year, calculate the after-tax net present value of this contract to Leslie, using a 6 percent discount rate. (Cash outflows and Negative amount should be indicated by a minus sign. Round discount factor(s) to 3 decimal places, intermediate calculations and final answers to the nearest whole dollar amount.) Year 0 Cash received Tax cost Net cash flow $ 0 Year 1: Cash received Tax cost Net cash flow Discount factor (6%) Present value of year 1 cash flow NPV Req A Req B>
Req A Req B Req C1 Req C2 French Corporation expects its marginal tax rate to be 31 percent both years. Calculate after-tax cost to enter into this contract using a 6 percent discount rate. (Cash outflows indicated by a minus sign. Round discount factor(s) to 3 decimal places, intermediate ca the nearest whole dollar amount.) Year 0 Cash paid Tax savings Net cash flow $ Year 1 Cash paid Tax savings Net cash flow $ Discount factor (6 % ) Present value of year 1 cash flow NPV Req A Req C1
Req C2 Req A Req B Req C1 Given that Leslie expects her tax rate to increase next year, she would prefer to receive more of the in project up front. Consider an alternative proposal under which French pays Leslie $48,000 this year, a year when the contract is complete. Calculate the after-tax benefit of this counterproposal to Leslie an to French. (Cash outflows and Negative amount should be indicated by a minus sign. Round discount f decimal places, intermediate calculations and final answers to the nearest whole dollar amount.) Value of restructured transaction to Leslie: Year 0 Cash received Tax cost Net cash flow $ 0 Year 1: Cash received Tax cost Net cash flow Discount factor (6%) Present value of year 1 cash flow NPV Cost of restructured transaction to French: Year 0 Cash paid Tax savings Net cash flow Year 1 Cash paid Tax savings Net cash flow $ Discount factor (6 %) Present value of year 1 cash flow NPV
Answer #1
LESLIE
Year0
A Cash received $34,000 (68000/2)
B=A*25% Tax Cost $8,500
C=A-B Net cash flow $25,500
Year1
D Cash received $34,000
E=D*35% Tax Cost $11,900
F=D-E Net cash flow $22,100
G Discount Factor (6%) 0.94339623 (1/(1+0.06)
H=F*G Present value of year1 cash flow $20,849
I=C+H NPV $46,349
FRENCH
Year0
A Cash paid $34,000
B=A*31% Tax Saving $10,540
C=A-B Net cash flow $23,460
Year1
D Cash paid $34,000
E=D*31% Tax Saving $10,540
F=D-E Net cash flow $23,460
G Discount Factor (6%) 0.94339623
H=F*G Present value of year1 cash flow $22,132
I=C+H NPV $45,592
VALUE OF RESTRUCTURED TRANSACTION TO LESLIE
Year0
Cash received $48,000
Tax Cost $12,000
Net cash flow $36,000
Year1
Cash received $18,000
Tax Cost $6,300
Net cash flow $11,700
Discount Factor (6%) 0.94339623
Present value of year1 cash flow $11,038
NPV $47,038
RESTRUCTURED TRANSACTION TO FRENCH
Year0
Cash paid $48,000
Tax Saving $14,880
Net cash flow $33,120
Year1
Cash paid $18,000
Tax Saving $5,580
Net cash flow $12,420
Discount Factor (6%) 0.94339623
Present value of year1 cash flow $11,717
NPV $44,837